OpenStudy (anonymous):

simplify cosec ( 2 sec^-1 3x + pi/2 ) in terms of x

OpenStudy (anonymous):

help ?

OpenStudy (jack1):

sorry, the jack has very little clue i can give u an answer, but not how i arrived at it?

OpenStudy (anonymous):

okay , at least you're helping

OpenStudy (jack1):

\[ \large \csc ( 2 sec^{-1} (3x) + \frac {\pi}{2}) \] \[ \large =\sec ( 2 sec^{-1} (3x)) \] \[ \huge = \frac {1}{ \frac {2}{9 x^2}-1}\] \[ \large = \frac {-9 x^2}{(9 x^2-2)}\] you'll have to ask @ganeshie8 or @hartnn to help out with the "why" if they're free ???

OpenStudy (anonymous):

@hartnn can you explain please ?

OpenStudy (anonymous):

by the way thanks jack

OpenStudy (kc_kennylau):

\[\csc\left(2\sec^{-1}(3x)+\frac\pi2\right)\]\[=\sec\left[\frac\pi2-\left(2\sec^{-1}(3x)+\frac\pi2\right)\right]\]\[=\sec\left(-2\sec^{-1}(3x)\right)\]\[=-\sec\left(2\sec^{-1}(3x)\right)\]\[=-\sec\left(2\sec^{-1}(3x)\right)\]

OpenStudy (anonymous):

can you explain it to me ?

OpenStudy (kc_kennylau):

\[=-\frac1{\cos\left(2\sec^{-1}(3x)\right)}\]\[=-\frac1{\cos^2\left(\sec^{-1}(3x)\right)-\sin^2\left(\sec^{-1}(3x)\right)}\] Let \(\theta\) be \(\sec^{-1}(3x)\), \(\sec\theta=3x\), \(\cos\theta=\dfrac1{3x}\). \[=-\frac1{\cos^2\theta-\sin^2\theta}\]\[=-\frac1{2\cos^2\theta-1}\]\[=-\frac1{\dfrac2{9x^2}-1}\]\[=-\frac1{\dfrac{2-9x^2}{9x^2}}\]\[=\frac{9x^2}{9x^2-2}\]

OpenStudy (kc_kennylau):

Because \(\sin\theta=\cos\left(\dfrac\pi2-\theta\right)\)

OpenStudy (kc_kennylau):

sorry gtg cya

OpenStudy (anonymous):

thanks , that's helping a lot !

OpenStudy (jack1):

that was masterfully done dude! @kc_kennylau props ;D

OpenStudy (kc_kennylau):

wow thanks for all the medals :D